OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

is it 2x or x^2

OpenStudy (jojo4eva):

11

OpenStudy (anonymous):

2^

OpenStudy (jojo4eva):

3^2=9+2=11

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

g(3)=11

OpenStudy (anonymous):

Thanks! @koook

OpenStudy (jojo4eva):

alrighty no love lol